Calculate the heat absorbed by the 100mL sample of water. The specific heat capacity of water is 4.18J/g℃.
1
Expert's answer
2020-10-02T14:19:59-0400

Specific heat capacity is the amount of heat required per gram of water to raise its temperature by "1\degree C" .

"C=4.18J/g \ \degree C"

Mass of water taken is "100 \ ml" and water density is "1g\/ml" .

So, "m=100\ g" .

If the water temperature rises from its initial temperature "T_1" to final temperature "T_2" , then change in temperature "\triangle T" "=(T_2-T_1)" .

Heat absorbed is given by "Q=mC\triangle T=mC(T_2-T_1)==100\times 4.18\triangle T\ J=418\triangle T \ J"

If temperature change is "\triangle T=50\degree C," then

"Q=418\times 50\ J=20900J=20.9\ KJ"